# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ 6. Technical Lemmas
In either case the inequality (2) follows from Lemma 6.2. Let $\mathfrak{S}$ denote the fan from the matching $\mathfrak{M}$ containing $v_{u_{n}}$ and $\mathfrak{S}^{\prime}$ denote $\mathfrak{S}$ with $v_{u_{n}}$ removed. The conditions of Lemma 6.2 are satisfied if we consider $\Gamma^{(n)}(-u_{n})$ as $\Gamma$ in that lemma, $v_{u_{n}}$ as $v$, and $\Gamma^{(n)\perp}(+\mathfrak{v})\cup\mathfrak{S}^{\prime}$ as $\Gamma^{\perp}$. Indeed, $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$ is $U^{(n)}(-u_{n},-v_{u_{n}})$-reflected. Moreover, the corresponding graph $\Gamma^{\star}$ from the lemma is obtained by removal of $\Gamma^{(n)\perp}(+\mathfrak{v})\cup\mathfrak{S}^{\prime}$ from $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$, therefore it is the same as $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ with $\mathfrak{S}$ removed. Since $\mathfrak{S}$ is a fan from a perfect $(1,d)$- matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$, we know that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition.
Therefore in case (i) by Lemma 6.2:
@@
[…]
Let us show that the inequality is strict. Indeed, by Lemma 6.2 the equality
@@
implies that $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(U_{X\setminus\{v_{u_{n}}\}})$ and $N_{\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})}(U_{X\setminus\{v_{u_{n}}\}})=\emptyset$. Therefore $v_{u_{n}}\in V^{(n)\perp}$. On the other hand, $v_{u_{n}}\in X\subset\Gamma^{(n)\star}$, a contradiction with the choice of $X$.
Again, inequalities $d\geq 2$ and $|X|>d$ implies
@@
Therefore, the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$ implies
@@
i.e. there exists some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$. Then the vertex $u_{j}^{\perp}$ belongs to $N_{\Gamma^{(n)}(-u_{n})\cap\mathcal{B}(u_{n})}(X)$ (recall that $X\subseteq\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$) and we find it effectively. Let us denote by $v^{\perp}_{j,1},\ldots,v^{\perp}_{j,d-1}$ the $(d-1)$-tuple of vertices adjacent to $u_{j}^{\perp}$ in $\Gamma^{(n)\perp}$.
@@
To finish the proof notice that the conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u_{n}$ is considered as $u$ and $u^{\perp}_{j}$ as $u^{\perp}$. Therefore by Lemma 6.6, $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem condition. ∎
@@
  • $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or $d-1$), are vertices adjacent to $u^{l+1}_{n}$ in $\mathfrak{M}^{2}_{n}$ (or in $\Gamma^{(n)\perp}$);
  • $\dot{v}_{1},\dot{v}_{2}$ are the elements among $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or $d-1$), which are not added to $M^{l+1}_{n}$; there are at most two of them.
@@
It is a counterpart of Claim 7.3. In the first case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}$ as $\Gamma^{\star}$ in Lemma 6.10, i.e. $\ell$ from this lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$. Thus $\ell\leq 3n+3$. Since $\mathfrak{T}$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}(+u_{j}^{\perp})$ as $\Gamma^{\star}$ i.e. $\ell$ from Lemma 6.10 is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$ and decreased by one, corresponding to the fan of $u_{j}^{\perp}$. Therefore it is at most $23n+$ 2. Since $\mathfrak{T}(+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. In the third case view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ as $\Gamma^{\star}$ i.e. $\ell$ from this lemma is equal to the number of fans in $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ increased by three, corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$ and decreased by two, corresponding to the fans of $u_{i}^{\perp},u_{j}^{\perp}$. Therefore it is at most $13n+$ 1. Since $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
@@
Assume that $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition. Let $u^{l+1}_{n}$ be the root of the last fan added to the matching $M_{n}$ in the second part of the $n$-th step; it belongs to the produced cycle of length $2$. Recall that $\dot{v}_{1},\dot{v}_{2}$ denote the vertices of the form $\dot{v}^{l+1}_{n,i}$ that were not added to $M^{l+1}_{n}$. In particular in Case 3A there is only one such a vertex and in Case 3B there is either one or two such vertices. From now on we consider the case of two additional vertices: $\dot{v}_{1}$ and $\dot{v}_{2}$. The case of only one of them, $\dot{v}_{1}$, is similar.
In the rest of the proof we will use the following two lemmas.
Lemma 7.8.
@@
We only consider the case when $\dot{v}_{2}$ exists, the other one is analogous. It follows from the procedure that:
@@
Since $\mathfrak{M}^{2}_{n}$ is a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$ and $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ is obtained from $\Gamma^{(n)\star}(-u_{n})$ by removal of two fans from $\mathfrak{M}^{2}_{n}$, by Proposition 6.3 we know that $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. ∎
Lemma 7.9.
@@
The inequality follows from Lemma 6.1. Indeed, consider $\Gamma^{(n+1)}$ to be $\Gamma$ from that lemma. Note that the construction guarantees that $\Gamma^{(n+1)}$ is $U^{(n+1)}$-reflected. Depending on existence of $\dot{v}_{2}$ consider either $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1})$ or $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1},+\dot{v}_{2})$ to be $\Gamma^{\perp}$ from that lemma. Then the corresponding graph $\Gamma^{\star}$ from the lemma is equal to either $\mathfrak{T}(-\dot{v}_{1})$ or $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ and by Lemma 7.8 it satisfies Hall’s $d$-harem condition. Therefore conditions of Lemma 6.1 are satisfied. ∎
Applying Lemma 7.8 we see that the graph $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. Let $\mathfrak{M}$ denote the perfect $(1,d)$-matching in $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$.
It is clear that the inequality $|N_{\mathfrak{T}}(X)|\geq|N_{\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})}(X)|$ holds for all $X\subset\mathfrak{U}$. This inequality also holds for subsets of $\mathfrak{V}$ which do not intersect $\{\dot{v}_{1},\dot{v}_{2}\}$. Therefore if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition then a witness of this is some finite subset of $\mathfrak{V}$ containing at least one of $\dot{v}_{1},\dot{v}_{2}$.
@@
Part 1. We check whether the graph $\mathfrak{T}(-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. If it does, we set
  • $M^{\prime}$ is a matching obtained by Proposition 5.1 in $\mathfrak{T}(-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$;
@@
  • $\Gamma^{\prime}=\mathfrak{T}$, denoting $\Gamma^{\prime}=(U^{\prime},V^{\prime})$,
@@
If it does not, then there exists a minimal connected set $X$ such that $\dot{v}_{1}\in X\subset\mathfrak{V}(-\dot{v}_{2})$ and $|N_{\mathfrak{T}(-\dot{v}_{2})}(X)|<\frac{1}{d}|X|$. Analogously to the Claim 7.4, it can be showed that such $X$ is a subset of $(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$ and as a consequence we can compute $X$ with these properties.
Observe that $|N_{\Gamma^{(n+1)}}(X)|\geq\frac{1}{d}|X|$ by Lemma 7.9.
@@
$M^{\prime}:=\mathfrak{M}^{2}_{n}\setminus\{(u^{l+1}_{n},\dot{v}^{l+1}_{n,1}),\ldots,(u^{l+1}_{n},\dot{v}^{l+1}_{n,d}),(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\}\cup$
@@
where $v^{\prime}_{n}=\dot{v}_{1}$. Observe that
@@
We have obtained $M^{\prime}$ from $\mathfrak{M}^{2}_{n}$ by removing the $d$-fan of $u^{l+1}_{n}$, adding the $d$-fan of $u^{\perp}_{j}$, and the following replacement: for each of $u^{\prime}_{i}$ we replace one edge incident to it by another incident edge (then $\dot{v}_{1}$ becomes adjacent to one edge in $M^{\prime}$). Since $\mathfrak{M}^{2}_{n}$ was a $(1,d)$-perfect matching in the ball $\mathcal{B}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$, it follows that matching $M^{\prime}$ satisfies the conditions of the perfect $(1,d)$-matchings for all susbsets of $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$. Therefore $M^{\prime}$ is a perfect $(1,d)$-matching in the ball $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$.
The conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$ and graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$ and $u^{\perp}_{j}$ is considered as $u^{\perp}$ in that lemma. Therefore by Lemma 6.6 graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition.
@@
Part 2. We check whether the graph $\Gamma^{\prime}$ satisfies Hall’s $d$-harem condition. If it does, then by Part 1 $\Gamma^{\prime}$ has to be equal to $\mathfrak{T}(+u_{j}^{\perp})$ and the proof is finished by the second option of the formulation.
If it does not then again, by repeating the reasoning of Part 1, we see that we can effectively recognize a minimal connected set $X$ such that $\dot{v}_{2}\in X\subset V^{\prime}\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$ and $|N_{\Gamma^{\prime}}(X)|<\frac{1}{d}|X|$. Again, using Lemma 7.9 we obtain
@@
i.e. there exists some $u^{\perp}_{i}\in N_{\Gamma^{(n+1)}}(X)$ and by high computability of $\Gamma^{(n+1)}$ it can be computed. As usual, we denote by $v^{\perp}_{i,1},\ldots,v^{\perp}_{i,d-1}$ the vertices adjacent to $u^{\perp}_{j}$ in $\dot{\Gamma}^{(n+1)\perp}$.
@@
$M^{\prime\prime}:=M^{\prime}\setminus\{(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\}\cup$
@@
We have obtained $M^{\prime\prime}$ from $M^{\prime}$ by adding the $d$ edges incident to $u^{\perp}_{i}$, and replacing one edge in matching for each of $u^{\prime}_{k}$ in such a way, that $\dot{v}_{2}$is adjacent to one edge in $M^{\prime\prime}$. It follows that matching $M^{\prime\prime}$ is the perfect $(1,d)$-matching in the ball $\Gamma^{\prime}(+u_{i}^{\perp})\cap(\mathcal{B}(u^{l+1}_{n})$.
The conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime\prime}$ as $\mathfrak{M}$ and $\Gamma^{\prime}(+u_{i}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$ and $u^{\perp}_{i}$ as $u^{\perp}$ in that lemma. Therefore by Lemma 6.6, $\Gamma^{\prime}(+u_{i}^{\perp})$ satisfies Hall’s $d$-harem condition.
The final argument depends on two possible outputs of Part 1. If the graph $\mathfrak{T}(-\dot{v}_{2})$ does not satisfy Hall’s $d$-harem condition (i.e. $u^{\perp}_{j}$ is involved), then $M^{\prime\prime}$ is a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$. If it does, we redefine $u^{\perp}_{j}:=u^{\perp}_{i}$ and then $M^{\prime\prime}$ becomes a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{j}^{\perp})$. Therefore if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition, then either $\mathfrak{T}(+u_{j}^{\perp})$ or $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies this condition. ∎

### v4

@@ 6. General Lemmas
We claim that in either case the inequality (2) follows from Lemma 6.2.
Figure 15. Application of Lemma 6.2; red letters correspond to the notation of Lemma 6.2.
Indeed, let $\mathfrak{S}$ denote the fan from the matching $\mathfrak{M}$ containing $v_{u_{n}}$ and $\mathfrak{S}^{\prime}$ denote $\mathfrak{S}$ with $v_{u_{n}}$ removed. The conditions of Lemma 6.2 are satisfied if we consider $\Gamma^{(n)}(-u_{n})$ as $\Gamma$ in that lemma, $v_{u_{n}}$ as $v$, and $\Gamma^{(n)\perp}\cup\{\mathfrak{v}\}\cup\mathfrak{S}^{\prime}$ as $\Gamma^{\perp}$. Indeed, $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$ is $U^{(n)}(-u_{n},-v_{u_{n}})$-reflected. Moreover, the corresponding graph $\Gamma^{\star}$ from Lemma 6.2 is obtained by removal of $\Gamma^{(n)\perp}\cup\{\mathfrak{v}\}\cup\mathfrak{S}^{\prime}$ from $\Gamma^{(n)}(-u_{n},-v_{u_{n}})$, therefore it is the same as $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$ with $\mathfrak{S}$ removed. Since $\mathfrak{S}$ is a fan from a perfect $(1,d)$-matching in $\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})$, we know that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition.
Therefore in case (i), by Lemma 6.2:
@@
 $\mbox{ i.e. }\,|N_{\Gamma^{(n)}(-u_{n})}(X)|\geq(d-1)(|X|-1)-1.$
Note that the inequality is strict. Indeed, by Lemma 6.2 the equality
@@
implies that $v_{u_{n}}\in N_{\Gamma^{(n)}(-u_{n})}(U_{X\setminus\{v_{u_{n}}\}})$ and $N_{\Gamma^{(n)\star}(-u_{n},-\mathfrak{v})}(U_{X\setminus\{v_{u_{n}}\}})=\emptyset$. Thus, $v_{u_{n}}\in V^{(n)\perp}$. On the other hand, $v_{u_{n}}\in X\subset\Gamma^{(n)\star}$, a contradiction with the choice of $X$.
Again, inequalities $d\geq 2$ and $|X|>d$ imply
@@
This finishes our argument for the statement that the assumption $|N_{\Gamma^{(n)\star}(-u_{n})}(X)|<\frac{1}{d}|X|$ implies
@@
i.e. there exists some $u_{j}^{\perp}\in N_{\Gamma^{(n)}(-u_{n})}(X)$.
The vertex $u_{j}^{\perp}$ that we have found, belongs to $N_{\Gamma^{(n)}(-u_{n})\cap\mathcal{B}(u_{n})}(X)$ (recall that $X\subseteq\mathcal{B}(u_{n})\setminus\mathcal{S}(u_{n})$). Thus, it can be found effectively by high computability of $\Gamma^{(n)}(-u_{n})$. Let us denote by $v^{\perp}_{j,1},\ldots,v^{\perp}_{j,d-1}$ the (d1)(d-1)-tuple of vertices adjacent to $u_{j}^{\perp}$ in $\Gamma^{(n)\perp}$.
$\bullet$ Step 2. $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem condition.
@@
To finish the proof, note that the conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$ and $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u_{n}$ is considered as $u$ and $u^{\perp}_{j}$ as $u^{\perp}$. Therefore, by Lemma 6.6, $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem condition. ∎
@@
  • $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or d1d-1), are vertices adjacent to $u^{l+1}_{n}$ in $\mathfrak{M}^{2}_{n}$ (or in $\Gamma^{(n)\perp}$);
  • $\dot{v}_{1},\dot{v}_{2}$ are the elements among $\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d$ (or d1d-1), which are not added to $M^{l+1}_{n}$; there are at most two of them.
@@
It is a counterpart of Claim 7.3. Apply Lemma 6.10. In the case of $\mathfrak{T}$, we view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}$ as $\Gamma^{\star}$ of that lemma. Furthermore, the number $\ell$ in its formulation is equal to the number of fans in both $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ and then increased by 3 (the number 3 corresponds to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$). Thus, $\ell\leq 3n+3$. Since $\mathfrak{T}$ satisfies Hall’s $d$-harem conditions, the lemma works.
In the second case we view $\Gamma^{(0)}$ as $\Gamma$ and $\mathfrak{T}(+u_{j}^{\perp})$ as $\Gamma^{\star}$ of Lemma 6.10. The number $\ell$ is equal to the number of fans in both $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ plus 3 (corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$) and minus 1 (corresponding to the fan of $u_{j}^{\perp}$). Therefore, it is at most $3n+2$. Since $\mathfrak{T}(+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again.
In the third case $\Gamma^{(0)}$ and $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ become $\Gamma$ and $\Gamma^{\star}$. The number $\ell$ from Lemma 6.10 is equal to the number of fans in both $\bigcup\limits_{i=1}^{n}M_{i-1}$ and $\Gamma^{(n)\perp}$ plus 3 (corresponding to the fans of $u_{n},u^{l+1}_{n}$ and $\dot{u}^{(n)\perp}$) and minus 2 (corresponding to the fans of $u_{i}^{\perp},u_{j}^{\perp}$). Therefore, $\ell\leq 3n+1$. Since $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies Hall’s $d$-harem conditions, Lemma 6.10 works again. ∎
@@
Assume that $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition. Let $u^{l+1}_{n}$ be the root of the last fan added to the matching $M_{n}$ in the second part of the $n$-th step; it belongs to the produced cycle of length $2$. Recall that $\dot{v}_{1},\dot{v}_{2}$ denote the vertices of the form $\dot{v}^{l+1}_{n,i}$ that were not added to $M^{l+1}_{n}$. In particular in Case 3A there is only one such a vertex and in Case 3B there is either one or two such vertices.
In the rest of the proof we will use the following two claims.
Claim 7.8.
@@
We only consider the case where $\dot{v}_{2}$ exists; the other case is analogous. It follows from the procedure of Section 5 that:
@@
Since $\mathfrak{M}^{2}_{n}$ is a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$ and $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ is obtained from $\Gamma^{(n)\star}(-u_{n})$ by removing two fans from $\mathfrak{M}^{2}_{n}$, by Proposition 6.3 we know that $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. ∎
Claim 7.9.
@@
The inequality follows from Lemma 6.1. Indeed, consider $\Gamma^{(n+1)}$ to be $\Gamma$ from that lemma. Note that the construction of Section 5 guarantees that $\Gamma^{(n+1)}$ is $U^{(n+1)}$-reflected. Depending on existence of $\dot{v}_{2}$, put either $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1})$ or $\dot{\Gamma}^{(n)\perp}(+\dot{v}_{1},+\dot{v}_{2})$ to be $\Gamma^{\perp}$ from that lemma. Then the corresponding graph $\Gamma^{\star}$ from the lemma is equal to either $\mathfrak{T}(-\dot{v}_{1})$ or $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ and by Claim 7.8 it satisfies Hall’s $d$-harem condition. Therefore, the conditions of Lemma 6.1 are satisfied. ∎
We now return to the main course of the proof. From now on we consider the case of two additional vertices: $\dot{v}_{1}$ and $\dot{v}_{2}$. The case of only one of them, $\dot{v}_{1}$, is similar and simpler (in particular, we do not need two parts in the further proof).
It is clear that the inequality $|N_{\mathfrak{T}}(X)|\geq|N_{\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})}(X)|$ holds for all $X\subset\mathfrak{U}$. This inequality also holds for subsets of $\mathfrak{V}$ which do not intersect $\{\dot{v}_{1},\dot{v}_{2}\}$. On the other hand, applying Claim 7.8 we see that the graph $\mathfrak{T}(-\dot{v}_{1},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. Therefore, if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition, then it would be witnessed by some finite subset of $\mathfrak{V}$ containing at least one of $\dot{v}_{1},\dot{v}_{2}$.
@@
$\bullet$ Part 1. In this part of the proof we define a graph $\Gamma^{\prime}$, a matching $M^{\prime}$ and a family of fans $\dot{\Gamma}^{(n+1)\perp}$. We start with checking whether the graph $\mathfrak{T}(-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition. If it does, we set
  • $M^{\prime}$ to be a matching obtained by Proposition 5.1 in $\mathfrak{T}(-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$;
@@
  • $\Gamma^{\prime}:=\mathfrak{T}$, denoting $\Gamma^{\prime}=(U^{\prime},V^{\prime})$,
@@
If the graph $\mathfrak{T}(-\dot{v}_{2})$ does not satisfy Hall’s $d$-harem condition, then there exists a minimal connected set $X$ such that $\dot{v}_{1}\in X\subset\mathfrak{V}(-\dot{v}_{2})$ and $|N_{\mathfrak{T}(-\dot{v}_{2})}(X)|<\frac{1}{d}|X|$. Applying the argument of Claim 7.4, one can show that such $X$ is a subset of $(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$. In particular, such a set $X$ can be computed.
Observe that $|N_{\Gamma^{(n+1)}}(X)|\geq\frac{1}{d}|X|$ by Claim 7.9.
@@
$M^{\prime}:=(\mathfrak{M}^{2}_{n}\setminus\{(u^{l+1}_{n},\dot{v}^{l+1}_{n,1}),\ldots,(u^{l+1}_{n},\dot{v}^{l+1}_{n,d}),(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\})\cup$
@@
where $v^{\prime}_{n}=\dot{v}_{1}$. We have obtained $M^{\prime}$ from $\mathfrak{M}^{2}_{n}$ by removing the $d$-fan of $u^{l+1}_{n}$, adding the $d$-fan of $u^{\perp}_{j}$, and the following replacement: for each of $u^{\prime}_{i}$ we replace one edge incident to it by another incident edge (then $\dot{v}_{1}$ becomes adjacent to one edge in $M^{\prime}$). Observe that
@@
Since $\mathfrak{M}^{2}_{n}$ was a $(1,d)$-perfect matching in the ball $\mathcal{B}(u^{l+1}_{n})\cap\Gamma^{(n)\star}$, the matching $M^{\prime}$ satisfies the conditions of the perfect $(1,d)$-matchings for all susbsets of $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$. Therefore, $M^{\prime}$ is a perfect $(1,d)$-matching in the ball $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})\cap\mathcal{B}(u^{l+1}_{n})$.
The conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime}$ as $\mathfrak{M}$, and the graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$, and $u^{\perp}_{j}$ is considered as $u^{\perp}$ in that lemma. Therefore, by Lemma 6.6 the graph $\mathfrak{T}(+u_{j}^{\perp},-\dot{v}_{2})$ satisfies Hall’s $d$-harem condition.
@@
$\bullet$ Part 2. We check whether the graph $\Gamma^{\prime}$ satisfies Hall’s $d$-harem condition. If it does, then by Part 1, $\Gamma^{\prime}$ has to be equal to $\mathfrak{T}(+u_{j}^{\perp})$ and the proof is finished by the second option of the formulation.
If the graph $\Gamma^{\prime}$ does not satisfy Hall’s $d$-harem condition, then repeating the reasoning of Part 1, we see that there is (and it can be effectively found) a minimal connected set $X$ such that $\dot{v}_{2}\in X\subset V^{\prime}\cap(\mathcal{B}(u^{l+1}_{n})\setminus\mathcal{S}(u^{l+1}_{n}))$ and $|N_{\Gamma^{\prime}}(X)|<\frac{1}{d}|X|$. Again, using Claim 7.9 we obtain
@@
i.e. there exists some $u^{\perp}_{i}\in N_{\Gamma^{(n+1)}}(X)$. By high computability of $\Gamma^{(n+1)}$, this element can be computed. As usual, we denote by $v^{\perp}_{i,1},\ldots,v^{\perp}_{i,d-1}$ the vertices adjacent to $u^{\perp}_{j}$ in $\dot{\Gamma}^{(n+1)\perp}$.
@@
$M^{\prime\prime}:=(M^{\prime}\setminus\{(u^{\prime}_{0},v^{\prime}_{0}),\ldots(u^{\prime}_{n-1},v^{\prime}_{n-1})\})\cup$
@@
We have obtained $M^{\prime\prime}$ from $M^{\prime}$ by adding $d$ edges incident to $u^{\perp}_{i}$, and replacing one edge in the matching for each of $u^{\prime}_{k}$ in such a way, that $\dot{v}_{2}$ is adjacent to an edge in $M^{\prime\prime}$. As a result, $M^{\prime\prime}$ is the perfect $(1,d)$-matching in the ball $\Gamma^{\prime}(+u_{i}^{\perp})\cap(\mathcal{B}(u^{l+1}_{n})$.
Now the conditions of Lemma 6.6 are satisfied if we consider $\Gamma^{(0)}$ as $\Gamma$ in this lemma, $\Gamma^{(n)\star}(-u_{n})$ as $\Gamma^{\star}$, $M^{\prime\prime}$ as $\mathfrak{M}$, and $\Gamma^{\prime}(+u_{i}^{\perp})$ as $(\Gamma^{\star}\setminus\mathfrak{F}_{u})\cup\mathfrak{S}$, i.e. $u^{l+1}_{n}$ is considered as $u$, and $u^{\perp}_{i}$ as $u^{\perp}$ in that lemma. Therefore, by Lemma 6.6, $\Gamma^{\prime}(+u_{i}^{\perp})$ satisfies Hall’s $d$-harem condition.
The final argument depends on two possible outputs of Part 1. If the graph $\mathfrak{T}(-\dot{v}_{2})$ does not satisfy Hall’s $d$-harem condition (i.e. $u^{\perp}_{j}$ is involved), then $M^{\prime\prime}$ is a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$. In the contrary case we redefine $u^{\perp}_{j}:=u^{\perp}_{i}$ and then $M^{\prime\prime}$ becomes a perfect $(1,d)$-matching in the graph $\mathfrak{T}(+u_{j}^{\perp})$. Therefore, if $\mathfrak{T}$ does not satisfy Hall’s $d$-harem condition, then either $\mathfrak{T}(+u_{j}^{\perp})$ or $\mathfrak{T}(+u_{i}^{\perp},+u_{j}^{\perp})$ satisfies this condition. ∎